a number cube is rolled 360 times and the results are 60ones, 58 twos, 70threes, 74 fours,38 fives, 60 sixes what are the probablity of rolling atwo or a three?
Answer by richard1234(7193)

Assuming the probabilities are independent, the probability of rolling a two or a three remains the same. This is equal to P(2) + P(3) = 1/6 + 1/6 = 1/3.